# Ultra-High-Energy Cosmic Rays from Neutrino-Emitting Tidal Disruption Events

### v2

@@ 2.2 UHECR injection in TDEs
We extend the model of Winter & Lunardini (2023) to nuclei following the methods in Boncioli et al. (2017); Biehl et al. (2018a, b), which we refer to for details. We simulate the source for different injection isotopes, which disintegrate and interact with the IR, OUV, and X-ray target photons in a test-particle approach, leading to a nuclear cascade in the source. Our computation includes the time evolution of the target photons, for which results will be discussed below. As main results, however, we present time-integrated neutrino and UHECR spectra since (at least for UHECRs) the time dependence cannot be measured. We then superimpose (sum) the results from different nuclei, which is an acceptable simplification since we do not take into account non-linear feedback (e.g., disintegration photons adding to the target).
@@
Our time-dependent solver is based on NeuCosmA, constructing a coupled Partial Differential Equation (PDE) system for all nuclear isotopes considered, protons, neutrons, muons, pions, kaons, and neutrinos, and evolving it with a Crank-Nicolson solver (Crank et al., 1947) of the PDE system re-parameterized on a double-log scale in -. The setup of the nuclear system is performed fully automatically with a recursive algorithm following all possible disintegration and beta decay channels beyond a chosen threshold value for the secondary multiplicity (1%) – which acts as a control parameter for the completeness of the system and for the precision of the results. Therefore, the size of the PDE system critically depends on the injection mass number , and the computation time scales correspondingly. Radiation processes taken into account are nuclear disintegration, spontaneous (nuclei) emissions, beta decays, synchrotron cooling losses, photo-pion production, Bethe Heitler pair production losses, escape (discussed below) and interaction losses, whereas we assume a stable production region without adiabatic losses. As a computational example a single run for AT2019fdr with 56Fe injection (largest considered) takes about two hours on a single workgroup server core. Out of 481 possible tabulated isotopes and 39962 tabulated disintegration channels (based on TALYS (Koning & Rochman, 2012)), 218 isotopes and 3838 disintegration channels are automatically selected in that case and evolved over the duration of the TDE. The same computation for proton injection takes only 30 seconds.
@@
For illustration, in this section, we follow the original models in Winter & Lunardini (2023) as closely as possible, which means that we use the same fiducial TDE parameters and also include an outflow with velocity , which the nuclei may interact with. Here we inject 56Fe with an ad hoc choice (whereas later will be varied). The relevant interaction and energy loss rates for the primary nucleus are shown in Fig. 1 for AT2019dsg. We see that, for 56Fe, nuclear disintegration dominates at the highest energies (thin blue curves), the source is optically thick, and the nuclear cascade will efficiently develop. Both the leading disintegration (thin blue) and photo-meson (thick blue, leading to neutrino production) exhibit a triple-hump structure from interactions off (from lowest to highest) the X-ray, OUV, and IR target photons. The solid curves refer to the peak time of the TDE, and the dashed curves refer to the time of the neutrino emission. Since the IR dust echo is delayed with respect to the blackbody peak, disintegration (and photo-meson production) due to scattering on the IR photons dominates at the time of the neutrino emission. Other processes are subleading at first but can become relevant for lighter nuclei over the evolution of the nuclear cascade as disintegration and photo-meson processes lead to nuclei with lower , see Biehl et al. (2018a) for details. Our simulation includes this evolution self-consistently in a fully time-dependent computation.
@@
It is an open question how UHECRs escape from the source. Here we define “escape spectra” as the cosmic ray emission spectra that are ejected into the circum-source environment. These escape spectra constitute the input of the UHECR propagation code, which describes transport through the cosmic microwave background and cosmic radiation fields. In the present scenario, CR nuclei at lower energies are confined in the production region by turbulent magnetic fields, whereas neutrons can free-stream out of the production region – unless it is optically thick to interactions. Since UHECRs can escape from the region for sufficiently large the Larmor radius, , a suitable parameterization is given by (Baerwald et al., 2013)
@@
with the diffusion coefficient in the Bohm-like regime and the free-streaming timescale . The first term ensures that the escape is limited by the free-streaming (ballistic) rate. While a different diffusion coefficient leads to a slightly different result, and also the coherence length of the magnetic field affects the shape of the escape component (see, e.g. Becker Tjus et al. (2022)), this approach produces relatively hard escape spectra, which are required by the UHECR data, see, e.g., Heinze et al. (2019). In Fig. 2, left panel, we show the corresponding cosmic-ray escape spectra, grouped into different mass groups (see figure caption) for AT2019dsg and pure 56Fe injection without transport effects (other than adiabatic losses). It is clear from the figure that the spectral shape below the peak of each component is quickly masked by a different component, which implies that, for hard enough escape spectra, the fit will not be very sensitive to the precise spectral index, and thus to the details of the diffusion model there. The H mass group is driven by protons and neutrons produced in the disintegration. The 56Fe (hypothetical) free-streaming escape spectrum is shown as the dashed curve, which corresponds to the in-source density suppressed by disintegration losses at the highest energies. It is lower than the blue Fe curve at the highest energies because it represents a single isotope only, and disintegration is very efficient, as can be seen from the shape of the curve and the from comparison with low energies, where disintegration does not occur. In fact, the escape is relatively inefficient compared to nuclear disintegration and photo-meson production, as can be seen in Fig. 1.
@@
There could be other contributions to UHECR escape. For example, the outflow may transport particles away by advection, which leads to softer escape spectra following the in-source density, see middle panel of Fig. 2. This contribution relies on the presence of the outflow and depends on its velocity. Another contribution would be the final release of all confined nuclei when the system ceases and magentic fields decay, which we illustrate in the right panel of Fig. 2. This contribution is, however, small because most of the energy of the system has been dissipated earlier. Since it turns out that the advective contribution cannot describe UHECR data very well, and the final release contribution is small, from here on we adopt the diffusive escape contribution only.222For self-consistency of the model, we nevertheless include the effects of the outflow in terms of interactions and an advective escape rate, which is similar to the effect of an adiabatic cooling rate – which may come from the expansion of the system. This leads to slightly smaller predicted neutrino fluences compared to Winter & Lunardini (2023) (where advective escape/adiabatic cooling was not considered) and to slightly harder proton spectra at low energies (where the protons cannot pile up so much anymore). While we only show the (summed) results over different mass groups in the figures in this section, we import the escape spectrum of each isotope individually into the UHECR transport code at the interface (see below).
@@
Even though we perform a fully time-dependent source simulation, for most of this study we will integrate the cosmic-ray escape spectra over time, because time-dependent effects for UHECRs are not measurable. However, the time-dependent escape is theoretically interesting, and the time-dependent UHECR spectra can have an effect on the time evolution of the neutrino luminosity. In Fig. 3 we show the time evolution of the diffusive UHECR escape luminosity for the different mass groups (colors) for AT2019dsg (left) and AT2019aalc (right), in comparison to the OUV and IR, and neutrino luminosities (black curves). The absolute luminosities should be interpreted with care because the UHECR luminosities are integrated in the energy range above (observer’s frame), and therefore effects at low energies are not captured. It is interesting that the primary nuclei disintegrate efficiently in the beginning, leading to the time-delayed (by the disintegration loss rate) appearance of lighter nuclei. There also can be a late-time rebrightening of the UHECR emission because the target photon fields decay and the injected UHECRs (following the OUV luminosity) can escape; this contribution is, however, small compared to the escape luminosity closer to the peak. Finally, note that, in the shown examples, the neutrino luminosities (dashed curves) peak close to the actual neutrino observation times better than for the proton model in Winter & Lunardini (2023). Here, the neutrino production is driven by the protons and neutrons at lower energies (see Fig. 2 and also Biehl et al. (2018a))), which are abundantly produced by the nuclear cascade and delayed by the disintegration cooling timescale.
@@ 2.3 Expected composition of accelerated cosmic rays from progenitors
In order to accurately reproduce the UHECR data, a significant portion of the accelerated nuclei must be predominantly composed of nitrogen and silicon groups, or, in other words, have a medium to medium-heavy composition (Aab et al., 2017; Heinze et al., 2019; Plotko et al., 2023). It is most natural to assume that these nuclei originate in the disrupted star and to identify their nuclear composition with the composition of the pre-disruption stellar matter. The latter is expected if there are no selection effects at play (favoring certain nuclear species over others) and if the black hole tidal forces do not change the composition of the stellar matter. Here, we consider several representative nuclear compositions motivated by the different types of disrupted stars. In particular, we examine main sequence stars (MS), ONeMg White Dwarfs (ONeMg-WD), CO White Dwarfs (CO-WD), Wolf-Rayet stars (WR), and red supergiants (RSG) (with composition averaged over their post-main-sequence lifetime), see Tab. 2. MS, RSG, and WR stars have similar mass abundances, primarily consisting of H and He with small amounts of medium and heavy elements. On the other hand, CO and ONeMg-WD stars have mass abundances dominated by medium and heavy elements, making them historically more suitable for UHECRs (Heinze et al., 2019; Aab et al., 2017).
In this study, we chose the injected nuclear species motivated by stellar evolution (LeBlanc, 2010): H, He, C, N, O, Na, Si, and Fe. Due to computation time restrictions and minimal impact on the final results, we couldn’t include all possible elements, so the fractions of elements not selected were added to the closest nuclear isotope. For the composition of the injected CR (), we consider two options. The first option involves a composition of injected CRs matching the mass abundance of the progenitor star (“Star Mass Abundance”). This assumption holds under specific conditions, such as having the same minimal Lorentz factor for all particles and a rigidity-dependent maximum energy (). In contrast, the second option treats the injected CR composition using free parameters, which are fit to data (“Fitted Composition Case”). The best-fit is then compared to the mass abundance of the chosen class of stars.
@@ 2.4 Consistency with individual neutrino-TDE associations
It’s important to validate the consistency of our individual TDE models with the observed data before studying the UHECRs from a population of TDEs. The key observation is the expected number of muon neutrinos for three specific TDEs: AT2019aalc, AT2019fdr, and AT2019fdsg (see Tab. 1). Our model involves three free parameters: the radius of the production region , the maximum rigidity , and the composition of injected CRs. It’s crucial to identify the allowed regions of these parameters that align with the expected number of muon neutrinos.
We start by calculating the muon neutrino fluence at Earth for each TDE across a grid for both the radius of the production region () and the maximum rigidity (), with the radius ranging from to cm and the maximum rigidity ranging from to GV. We consider pure injection for CRs of a single isotope from the seven possible injection isotopes (H, He, C, N, O, Na, Si, Fe). The next ingredient is effective area , which represents the IceCube detector’s sensitivity to neutrinos across different energy ranges. We utilize the Gamma-Ray Follow-up (GFU) effective area as described in (Blaufuss et al., 2019). The predicted number of muon neutrino events by a model is then calculated by combining the predicted fluences with the effective areas over the entire energy range. This calculation can be expressed as:
The next step involves comparing the predicted number of muon neutrinos expected by IceCube. Based on the previous studies, there are certain expectations for the number of muon neutrinos from AT2019dsg and AT2019fdr, but such expectations are not available for AT2019aalc. For instance, the detection of a single high-energy neutrino associated with AT2019dsg suggests a mean expectation in the range of 0.008 to 0.76 muon neutrinos at 90% confidence, considering the total TDE population observed by the Zwicky Transient Facility (ZTF) (Stein et al., 2021). Similarly, for AT2019fdr, the expected number of muon neutrinos is between 0.007 and 0.13, based on its contribution to the g-band peak energy flux among nuclear transients (Reusch et al., 2022). Based on this information, we use the allowed range of 0.007-0.13 muon neutrinos for our study. As indicated in the study by Winter & Lunardini (2023), the radius of the production region is the main control parameter for the expected number of neutrinos. Consequently, we present our findings graphically against the radius as shown in Fig. 4. For each value of the radius of the production region, we calculate the minimum and maximum numbers of neutrinos across the entire range of maximum rigidity and all injected elements.
@@
In terms of the allowed region for the radius, the expected number of neutrinos for AT2019fdr and AT2019dsg, shown as the grey-blue and red curve in Fig. 4, falls within the allowed neutrino region (marked as a grey area on the plot) only at smaller radii. AT2019dsg shows agreement with expected values only at the smallest radii from to cm. AT2019fdr indicates a preference for smaller radii falling within the range of to cm. Since AT2019aalc is the most efficient neutrino TDE, the allowed region is shifted to a bigger radius value from to cm. Besides the restriction from the expected number of muon neutrinos, there is another independent method to estimate upper limits on the production radius involving IR observations. By interpreting the time-delayed IR emission as a dust echo, the size of the production region can be estimated (see in Tab. 1, and corresponding references). For all TDEs, the allowed region based on the expected number of neutrinos is inside the one based on IR data (grey arrows on the bottom plot in Fig. 4).
Apart from influencing the expected number of neutrinos, the production radius also significantly affects the CR disintegration within the radiation zone. Fig. 5 shows the escaped spectra of CRs and neutrinos for two distinct radii — and cm — while maintaining the same maximum rigidity ( GV) and CR injection (pure Si). A smaller radius increases the CR interactions with background photons before they can escape, leading to a higher flux of secondary particles at similar energies of the injected primary element. Conversely, a larger radius allows CR survival. As demonstrated by Aab et al. (2017) and Heinze et al. (2019), UHECR data indicate sources with a peaky quasi-monochromatic spectrum of isotopes transitioning from protons at lower energies to iron at higher energies. For our study, this means that a “contamination” of secondary nuclei in the escaped spectrum from disintegration makes the source unsuitable as UHECR source. This especially affects AT2019dsg, for which fixing the radius to smaller values, as required to produce a high enough neutrino fluence, implies such disintegration. Because of this tension, we assume that the dsg-like TDEs do not significantly contribute to UHECR and exclude it from our analyses. The following section will focus only on AT2019aalc and AT2019fdr.

### v3

@@ 2.2 UHECR injection in TDEs
We extend the model of W. Winter & C. Lunardini (2023) to nuclei following the methods in D. Boncioli et al. (2017); D. Biehl et al. (2018a, b), which we refer to for details. We simulate the source for different injection isotopes, which disintegrate and interact with the IR, OUV, and X-ray target photons in a test-particle approach, leading to a nuclear cascade in the source. We perform the computation, including the full-time evolution of the target photons, which we also show below; for the main results, we use time-integrated neutrino and UHECR spectra since (at least for UHECRs) the time dependence cannot be measured. We then superimpose the results from different nuclei in the source-propagation chain, which is a possible simplification, since we do not take into account non-linear feedback (e.g., disintegration photons adding to the target).
@@
Our time-dependent solver is based on NeuCosmA, constructing a coupled Partial Differential Equation (PDE) system for all nuclear isotopes considered, protons, neutrons, muons, pions, kaons, and neutrinos, and evolving it with a Crank-Nicolson solver (J. Crank et al., 1947) of the PDE system re-parameterized on a double-log scale in -. The setup of the nuclear system is performed fully automatically with a recursive algorithm following all possible disintegration and beta decay channels beyond a chosen threshold value for the secondary multiplicity (1%) – which acts as a control parameter for the completeness of the system and for the precision of the results. Therefore, the size of the PDE system critically depends on the injection mass number , and the computation time scales correspondingly. Radiation processes taken into account are nuclear disintegration, spontaneous (nuclei) emissions, beta decays, synchrotron cooling losses, photo-pion production, Bethe Heitler pair production losses, escape (discussed below) and interaction losses, whereas we assume a stable production region (), which means that adiabatic losses can be neglected. As a computational example a single run for AT2019fdr with 56Fe injection (largest considered) takes about two hours on a single workgroup server core. Out of 481 possible tabulated isotopes and 39962 tabulated disintegration channels (based on TALYS (A. J. Koning & D. Rochman, 2012)), 218 isotopes and 3838 disintegration channels are automatically selected in that case and evolved over the duration of the TDE. The same computation for proton injection takes only 30 seconds.
corresponds to the evolution of the whole system, and the free-streaming rate is . The shaded region is beyond the assumed maximal injection energy.
@@
For illustration in this section, we follow the original models in W. Winter & C. Lunardini (2023) as closely as possible, which means that we use the same fiducial TDE parameters and also include an outflow with the nuclei may interact with. Here we inject 56Fe with an ad hoc choice (whereas later will be varied). The relevant interaction and energy loss rates for the primary nucleus are shown in Fig. 1 for AT2019dsg. What we see here is that for iron, nuclear disintegration dominates at the highest energies (thin blue curves), the source is optically thick, and the nuclear cascade will efficiently develop. Both the leading disintegration (thin blue) and photo-meson (thick blue, leading to neutrino production) exhibit a triple-hump structure from interactions off (from lowest to highest) the X-ray, OUV, and IR target photons. Note that the solid curves refer to the peak time () of the TDE, and dashed curves refer to the time of the neutrino emission; since the IR dust echo is delayed with respect to the blackbody peak, disintegration (and photo-meson production) of the IR photons dominates at the time of the neutrino emission. Other processes are subleading at first but can become relevant for lighter nuclei over the evolution of the nuclear cascade as disintegration and photo-meson processes lead to nuclei with lower , see D. Biehl et al. (2018a) for details. Our simulation includes this evolution self-consistently in a fully time-dependent computation.
@@
It is an open question how UHECRs escape from the source, where “escape spectra” refer to the cosmic ray emission spectra ejected into the circum-source environment and injected into the UHECR propagation code, which describes transport through cosmic microwave background and cosmic radiation fields. In the present scenario, CRs at lower energies are confined in the production region by turbulent magnetic fields, whereas the electrically neutral neutrons can free-stream out of the production region – unless it is optically thick to interactions. Since UHECRs can escape from the region if the Larmor radius , a suitable parameterization is given by (P. Baerwald et al., 2013)
@@
with the diffusion coefficient in the Bohm-like regime and the free-streaming timescale . The first term ensures that the escape is limited by the free-streaming (ballistic) rate. While a different diffusion coefficient leads to a slightly different result, and also the coherence length of the magnetic field affects the shape of the escape component (see, e.g. J. Becker Tjus et al. (2022)), this approach describes relatively hard escape spectra required from UHECR data, see, e.g., J. Heinze et al. (2019). We show in Fig. 2, left panel, the corresponding cosmic-ray escape spectra group into different mass groups (see figure caption) for AT2019dsg and pure 56Fe injection without transport effects (other than adiabatic losses). It is clear from the figure that the spectral shape below the peak of each component is quickly masked by a different component, which implies that, for hard enough escape spectra, the fit will not be very sensitive to the precise spectral index, and thus to the details of the diffusion model there. The H mass group is driven by protons and neutrons produced in the disintegration, which is suppressed by photo-meson production at high energies. The 56Fe (hypothetical) free-streaming escape spectrum is shown as the dashed curve, which corresponds to the in-source density suppressed by disintegration losses. It is lower than the Fe curve because it represents a single isotope only, and disintegration is very efficient, as can be seen from the shape of the curve and the comparison with low energies, where disintegration does not occur. In fact, the escape is relatively inefficient compared to nuclear disintegration and photo-meson production, as can be seen in Fig. 1.
@@
One could also think about other contributions to UHECR escape. For example, the outflow may transport particles away by advection, which leads to softer escape spectra following the in-source density, see middle panel of Fig. 2. This contribution relies on the presence of the outflow and depends on its velocity. Another contribution would be the final release of all confined nuclei when the system (such as magnetic fields) decay, which we illustrate in the right panel of Fig. 2. This contribution is, however, small because most of the energy of the system has been dissipated otherwise. Since it turns out that the advective contribution cannot describe UHECR data very well, and the final release contribution is small, we further adopt the diffusive contribution only.333For self-consistency, we nevertheless include the effects of the outflow in terms of interactions and an advective escape rate, which is similar to the effect of an adiabatic cooling rate. This leads to slightly smaller predicted neutrino fluences compared to W. Winter & C. Lunardini (2023) (where advective escape/adiabatic cooling was not considered) and to slightly harder proton spectra at low energies (where the protons cannot pile up so much anymore). While we only show the (summed) results over different mass groups in the figures in this section, we transfer all escape spectra of all isotopes in the system to the UHECR transport code at the interface (see below).
@@
Even though we perform a fully time-dependent source simulation, we will, for most of this study, integrate the escaping cosmic-ray spectra over time because time-dependent effects for UHECRs are not measurable. However, the time-dependent escape is theoretically interesting, and the time-dependent spectra can have an effect on the neutrino production time. We, therefore, show in Fig. 3 the time evolution of the diffusive UHECR escape luminosity for the different mass groups (colors) for AT2019dsg (left) and AT2019aalc (right), in comparison to the OUV and IR, and neutrino luminosities (black curves). The absolute luminosities should be interpreted with care because the UHECR luminosities are integrated in the energy range above (observer’s frame), and therefore effects at low energies are not captured. It is interesting that the primary nuclei disintegrate efficiently in the beginning, leading to the time-delayed (by the disintegration loss rate) appearance of lighter nuclei. There also can be a late-time recovery of UHECR escape because the target photon fields decay and the injected UHECRs (following the OUV luminosity) can escape; this contribution is, however, small compared to the escape luminosity closer to the peak. Finally, note that, in the shown examples, the neutrino luminosities (dashed curves) peak close to the actual neutrino observation times better than for the proton model in W. Winter & C. Lunardini (2023). Here, the neutrino production is driven by the protons and neutrons at lower energies (see Fig. 2 and also D. Biehl et al. (2018a)), which are abundantly produced by the nuclear cascade and delayed by the disintegration cooling timescale.
@@ 2.3 Expected composition of accelerated cosmic rays from progenitors
In order to accurately replicate UHECR data, a significant portion of the accelerated nuclei must be predominantly from the nitrogen and silicon groups, i.e., have a medium to medium-heavy composition (A. Aab et al., 2017; J. Heinze et al., 2019; P. Plotko et al., 2023). It is most natural to assume that these nuclei originate from the disrupted star and to identify their nuclear composition with the composition of the progenitor. That direct relationship is expected if there are no selection effects at play (e.g., favoring the acceleration of certain nuclear species over others) and if the black hole tidal forces do not change the composition of the stellar matter. Here, we consider several representative nuclear compositions motivated by the different types of disrupted stars. In particular, we examine main sequence stars (MS), ONeMg White Dwarfs (ONeMg-WD), CO White Dwarfs (CO-WD), Wolf-Rayet stars (WR), and red supergiants (RSG) (with composition averaged over their post-main-sequence lifetime), see Tab. 2. MS, RSG, and WR stars have similar mass abundances, primarily consisting of H and He with small amounts of medium and heavy elements. On the other hand, CO and ONeMg-WD stars have mass abundances dominated by medium and heavy elements, making them historically more suitable for UHECRs.
In this study, we chose the injected nuclear composition motivated by stellar evolution (F. LeBlanc, 2010): H, He, C, N, O, Na, Si, and Fe. Due to computation time restrictions and minimal impact on the final results, we couldn’t include all possible elements, so the fractions of elements not selected were added to the closest nuclear isotope. For the composition of the injected CR (), we consider two options. The first option involves a composition of injected CRs matching the mass abundance. This assumption holds under specific conditions, such as having the same minimal Lorentz factor for all particles and a rigidity-dependent maximum energy (). In contrast, the second option treats the injected CR composition as a free parameter and compares it to the mass abundance of a chosen set of stars.
@@
2.4 Tidal disruption constraints on SMBH mass from different progenitors
A star is tidally disrupted when its tidal disruption radius exceeds the Schwarzschild radius of the black hole; here is the SMBH mass, is the star’s radius, is the star’s mass and is the gravitational constant. If , the star will cross the event horizon before being torn apart, preventing the formation of a luminous TDE. The maximum SMBH mass for which a star of given mass and radius can be disrupted is given by S. Rosswog et al. (2009):
Description | Symbol | Star’s stage | White Dwarf | |||
MS | RSG | WR | CO-WD | ONeMg-WD | ||
Stellar mass [] | 1.0 | 7.5 | 0.6 | 1.1 | ||
Stellar radius [] | 1.0 | 3.6 | ||||
Upper limit on SMBH mass [] |
Tab. 3 lists the upper limits on SMBH mass for different progenitor types. The SMBHs associated with AT2019dsg, AT2019fdr, and AT2019aalc have estimated masses around , indicating that main sequence stars, red supergiants, and Wolf-Rayet stars can be tidally disrupted in these events. However, white dwarfs (both CO-WD and ONeMg-WD) have much lower values, suggesting that they would likely plunge directly into the black hole before experiencing tidal disruption.
These calculations assume a non-rotating (Schwarzschild) black hole, whereas for of a spinning (Kerr) black hole, the disruption conditions are modified. A rapidly rotating SMBH increases the tidal disruption radius relative to the event horizon, effectively increasing the maximum mass for which disruption can occur. For a maximally spinning Kerr black hole, can be up to 1.5-10 times higher than in the non-spinning case (H. Sponholz, 1994; S. Kobayashi et al., 2004; A. Mummery, 2023). This means that white dwarfs that would otherwise be swallowed as a whole by a non-rotating SMBH could still undergo partial disruption near a rapidly spinning SMBH.
While the direct disruption of white dwarfs in AT2019dsg, AT2019fdr, and AT2019aalc appears highly unlikely, even when considering spinning effects and uncertainties in the SMBH masses associated with these events, we include WD-like compositions in our modeling for comparison. This is motivated by previous studies (B. T. Zhang et al., 2017; D. Biehl et al., 2018b), which indicate that a white dwarf-like composition are necessary to explain the observed cosmic-ray composition at Earth. Although our primary focus is on fitting the composition to the available data, with main-sequence stars as the dominant progenitors, it is important to emphasize this clear tension between the high SMBH masses, which strongly disfavor white dwarf disruptions in these events, and the heavy-element composition inferred from UHECR observation.
@@ 2.5 Consistency with individual neutrino-TDE associations
It’s important to validate the consistency of our individual TDE models with the observed data before studying the UHECRs from a population of TDEs. The key observation is the expected number of muon neutrinos for three specific TDEs/candidates: AT2019aalc, AT2019fdr, and AT2019fdsg (see Tab. 1). Our model involves three free parameters: the radius of the production region , the maximum rigidity , and the composition of injected CRs. It’s crucial to identify the allowed regions of these parameters that align with the expected number of muon neutrinos.
We start by calculating the muon neutrino fluence at Earth for each TDE across a grid for both the radius of the production region () and the maximum rigidity (), with the radius ranging from to cm and the maximum rigidity ranging from to GV. We consider pure injection for CRs of a single isotope from the seven possible injection isotopes (H, He, C, N, O, Na, Si, Fe). The next ingredient is effective area , which represents the IceCube detector’s sensitivity to neutrinos across different energy ranges. We utilize the Gamma-Ray Follow-up (GFU) effective area as described in (E. Blaufuss et al., 2019). The predicted number of muon neutrino events by a model is then calculated by combining the predicted fluences with the effective areas over the entire energy range. This calculation can be expressed as:
The next step involves comparing the predicted number of muon neutrinos expected by IceCube. Based on the previous studies, there are certain expectations for the number of muon neutrinos from AT2019dsg and AT2019fdr, but such expectations are not available for AT2019aalc. For instance, the detection of a single high-energy neutrino associated with AT2019dsg suggests a mean expectation in the range of 0.008 to 0.76 muon neutrinos at the 90% confidence level, considering the total TDE population observed by the Zwicky Transient Facility (ZTF) (R. Stein et al., 2021). Similarly, for AT2019fdr, the expected number of muon neutrinos is between 0.007 and 0.13, based on its contribution to the g-band peak energy flux among nuclear transients (S. Reusch et al., 2022). Based on this information, we use the allowed range of 0.007-0.13 muon neutrinos for our study. As indicated in the study by W. Winter & C. Lunardini (2023), the radius of the production region is the main control parameter for the expected number of neutrinos. Consequently, we present our findings graphically against the radius as shown in Fig. 4. For each value of the radius of the production region, we calculate the minimum and maximum numbers of neutrinos across the entire range of maximum rigidity and all injected elements.
@@
In terms of the allowed region for the radius, the expected numbers of neutrinos for AT2019fdr and AT2019dsg, shown as the grey-blue and red curves in Fig. 4, fall within the allowed neutrino region (marked as a grey area on the plot) only at smaller radii. AT2019dsg shows agreement with expected values only at the smallest radii from to cm. AT2019fdr indicates a preference for smaller radii falling within the range of to cm. Since AT2019aalc is the most efficient neutrino TDE, the allowed region is shifted to a bigger radius value from to cm. Besides the restriction from the expected number of muon neutrinos, there is another independent method to estimate upper limits on the production radius involving IR observations. By interpreting the time-delayed IR emission as a dust echo, the size of the production region can be estimated (see Tab. 1 and corresponding references). For all TDEs, the allowed region based on the expected number of neutrinos is inside the one based on IR data (grey arrows on the bottom plot in Fig. 4).
Apart from influencing the expected number of neutrinos, the production radius also significantly affects the CR disintegration within the radiation zone. Fig. 5 shows the escaped spectra of CRs and neutrinos for two distinct radii — and cm — while maintaining the same maximum rigidity ( GV) and CR injection (pure Si). A smaller radius increases the CR interactions with background photons before they can escape, leading to a higher flux of secondary particles at similar energies of the injected primary element. Conversely, a larger radius allows CR survival. As demonstrated by A. Aab et al. (2017) and J. Heinze et al. (2019), UHECR data indicate sources with a peaky quasi-monochromatic spectrum of isotopes transitioning from protons at lower energies to iron at higher energies. For our study, this means that a “contamination” of secondary nuclei in the escaped spectrum from disintegration makes the source unsuitable as UHECR source. This especially affects AT2019dsg, for which fixing the radius to smaller values, as required to produce a high enough neutrino fluence, implies such disintegration. Consequently, we assume that the dsg-like TDEs do not significantly contribute to UHECR and exclude it from our analyses. The following section will focus only on AT2019aalc and AT2019fdr.